# Designing and Evaluating Hint Generation Systems for Science Education

### v1

@@ 5.3. Static vs Dynamic hints
We observed minimal differences in participant performance across the static and dynamic hint conditions, with average success of 7.61/10 and 7.88/10 questions respectively, compared to 6.36/10 questions in the no-hint control setting. Participants requested a total of 639 static hints, and 510 dynamic hints. The perceived difficulty of quiz was also rated similarly across all three sections, suggesting that neither hint strategy provided a decisive advantage at the level of overall outcomes. A Mann-Whitney U test (Mann and Whitney, 1947) revealed no statistically significant difference between the satisfaction ratings ($U=168351,p=0.365$) with negligible effect size (rank-biserial correlation = -0.033). Even after revealing the corresponding hint strategy to participants in the post-quiz survey, opinions remained divided. 20 out of 41 participants found static hints to be more helpful for answering the questions, while 23 out of 41 participants felt dynamic hints better improved their knowledge about the questions. Point-biserial correlation (Kornbrot, 2014) between post-section responses and post-quiz responses revealed that the reported quiz difficulty was not correlated with the post-disclosure preference ($corr=0.067,p=0.675$), indicating that the participants independently rated the quiz difficulty agnostic of hint helpfulness. The post-section hint helpfulness was moderately correlated with the post disclosure informativeness preference obtained from the posttest survey ($corr=0.471,p=0.002$), indicating that participants who found one hint strategy more helpful were likely to maintain their preference even after knowing the hint strategy. However, there is no correlation between the performance difference between static and dynamic, and the post-disclosure preference of answering assistance ($corr=0.063,p=0.694$), indicating that participants’ preferences were not influenced by their actual performance but rather by their subjective experience with the hints. Although neither strategy prevailed overall, participants described clear differences in how they experienced the two hinting approaches.
@@
Static hints–pre-generated in chains of four–were rigid in their progression but often provided diverse facts that enriched the context of the question. For example, ”Reflect on the types of landforms created by magma rising to the surface.”, or ”Research the type of radiation that includes visible light.” (see Figure 5). This pattern was reflected in reported informativeness scores: 39% of static hint chains were labeled informative, compared to only 35% of dynamic chains. However, Chi-squared test (Tallarida and Murray, 1987) revealed no statistically significant difference between the two strategies ($\chi^{2}=0.819,p=0.356$) with no association (Cramer’s V = 0.027). Providing a breadth of information allowed static hints to support sense-making and knowledge-building, but at times the progression felt misaligned with participants’ immediate needs.
@@
Dynamic hints, in contrast, were answer-fixated and more directly oriented toward guiding participants to the correct response. Because these hints were generated adaptively without awareness of the total number of hints in the chain, they lacked the deliberate pacing and planning of the static setting. However, this immediacy was effective when participants lacked sufficient prior knowledge to answer the question. For instance, hints like, ”Think about the rays named after a Greek letter”, or ”Consider a medication that starts with ”a” and is often used for headaches.” (see Figure 5). At the same time, the absence of chain-level planning made it harder to conceal answers: leakage was reported more often in dynamic hints (18%) than in static hints (15%). While Chi-squared test (Tallarida and Murray, 1987) revealed a statistically significance difference between the leakage responses ($\chi^{2}=5.35,p=0.021$), there was no association (Cramer’s V = 0.068). If the participant’s did not engage with the question, (i.e., not attempting any answers) the average satisfaction ratings for dynamic hints was worse than the static hints without engagement. While when they engaged with dynamic hints, their satisfaction ratings were comparable to static hints (see Figure 6 bottom).
@@
Across participants, the overall sentiment remained consistent. Static hints were valued for their contextual richness, while dynamic hints were appreciated for their answer-directed efficiency. Minor exceptions arose due to randomized question distribution, with four participants noting one section to be easier than other, but the broader trend was stable. These findings suggest that static hints may be better suited for fostering understanding and providing transferable knowledge, whereas dynamic hints may be more effective in ensuring task success when learners are struggling.
@@ 5.4. Assessing effectiveness of automatic evaluation metrics
Automatic evaluation metrics are a key step toward scaling hint generation research: they allow us to compare systems without relying exclusively on human annotation. However, their usefulness depends on how well they capture the aspects of hint quality that learners themselves experience. To assess this alignment, we compared a set of automatic evaluation metrics against participant feedback collected in our human study. In total, 41 participants annotated 1,149 hints across 468 questions. Unlike our standardized automatic evaluation setup with a fixed chain size (see Section A), the hint chains here varied in length, providing a more realistic but also more heterogeneous basis for evaluation.
@@
We observed only weak correlations between information gain metrics and participants’ reported informativeness of hints for $InfoGain_{mean}$ ($corr_{pearson}=-0.121$) and $InfoGain_{comb}$ ($corr_{pearson}=-0.142$). These negative but shallow correlations highlight the limitations of relying on LLM-based measures of content quality, which often fail to capture the contextual insightfulness valued by learners. Although our study design did not directly test $Redundancy$ or $Consistency$, we also examined their correlations with user satisfaction. $Redundancy$ showed a weak negative correlation ($corr_{pearson}=-0.245$), while $Consistency$ was essentially uncorrelated ($corr_{pearson}=-0.053$).
@@
For answer leakage, we were able to compare participant reports directly to our automatic metrics. The exact-match metric achieved high precision (0.834) but very low recall (0.075), indicating that it severely underestimated the degree of leakage present. While we implemented $leakage_{EM}$ with a simple string matching approach, that is unable to capture the nuances where hints reveal the answer in plural form (e.g., capillary vs capillaries), or change in part-of-speech (e.g., volcano vs volcanic). We also identified several indirect oversimplifications from hints that remain undetected by a simple syntactic overlap approach. For example, ”The bohr model works only for which atom?”, participants considered the hint ”It is the first element on the periodic table.” to have given away the answer, whereas it was undetected by the $leakage_{EM}$ metric. Conversely, the LLM-based leakage metric offered slightly higher recall (0.183) but suffered from low precision (0.266), indicating that while the learners didn’t consider certain hints to be oversimplifications, the evaluation LLM considered those hint chains to be so, overshooting the answer leakage predictions. For example, for the question ”What is the most abundant metal of the earth’s crust?”, and the provided hints ”Consider which metals are commonly found in rocks and soil.” and ”Think about the primary component of bauxite, a common ore.”, $leakage_{LLM}$ metric considers the hint chain to have leaked the answer (Aluminum), whereas the participant doesn’t indicate that in their responses. These results suggest that neither approach yet provides a reliable estimate of answer leakage from a learner’s perspective.
@@ 6. Discussion
Our findings highlight the complexity of automatic hint generation and the limitations of adopting off-the-shelf LLMs as tutoring systems. While participants consistently valued certain hint characteristics such as simplicity, novel knowledge perspectives, content diversity, and incremental scaffolding, their preferences diverged on various other aspects. For instance, some learners favored answer-focused hints that guided them toward the answer, while others found such feedback unhelpful in improving their learning. Beyond feedback on the hints generated by our system, participants also emphasized some unmet needs, with the hints unable to help bridge the linguistic barriers or connect well to their prior knowledge. Below, we discuss design considerations for advancing hint generation beyond factual question answering (Section 6.1), and we position this research problem as a potential avenue for addressing the growing challenge of cognitive offloading among learners (Section 6.2).
@@ 6.1. Design Implications Beyond Factual Hinting
The revised Bloom’s taxonomy of educational goals (Bloom et al., 1956; Anderson and Krathwohl, 2001) organizes cognitive processes from remember and understand through apply, analyze, evaluate, and create, while distinguishing knowledge goals between factual, conceptual, procedural, and metacognitive dimensions. Our study deliberately focused on factual knowledge at the lower levels of this hierarchy, where hints primarily supported remembering and understanding. This scope allowed us to systematically examine the role of hinting, but also revealed several limitations in current systems, for instance, challenges in robust answer assessment and misalignment between automatic evaluation metrics and learner perceptions. As one moves upward in Bloom’s hierarchy, such limitations become even more consequential: supporting conceptual reasoning, procedural mastery, or metacognitive reflection requires richer forms of adaptation, sensitivity to learner goals, and integration of affective and self-regulatory scaffolding. Next, we outline design implications that extend beyond factual hinting, drawing on our findings to point toward future directions for building tutoring systems capable of supporting higher-order learning.
6.1.1. Building Trustworthy Foundations
@@
Automatic Evaluation of Hint Quality. Our study also highlights the limitations of existing automatic evaluation metrics for hint generation. While such metrics are essential for scaling research without relying exclusively on human annotation, our analysis revealed that they are, at best, weakly aligned with learners’ perceptions of hint quality. For example, information gain metrics showed shallow negative correlations with participant ratings of informativeness, suggesting that LLM–based measures fail to capture the contextual insightfulness learners value. Similarly, our automatic leakage metrics exhibited trade-off. Exact-match rules achieved high precision but severely underestimated leakage, while LLM-based approaches captured more cases but at the cost of low precision. Redundancy and consistency measures also showed little to no correlation with learner satisfaction. These findings suggest that current metrics do not yet provide a reliable proxy for learner experience. Progress in this space will require either refining existing metrics or developing new ones grounded in interaction data and learner-centered perspectives. Our annotated dataset of 1100+ hints across 450+ questions could support development of more trustworthy evaluation frameworks.
@@
LLM-as-a-Judge. Using LLMs as evaluators of natural language output-commonly referred to as LLM-as-a-judge-has gained traction in NLP research (Gu et al., 2024). Compared to resource-intensive human studies, it offers easy setup, low expense, and circumvents challenges such as IRB approvals, funding constraints, and participant recruitment. Yet this approach comes with significant limitations. Prior work shows that LLM-based evaluators are prone to systematic biases (Chen et al., 2024a; Ye et al., 2024; Shi et al., 2024; Wang et al., 2023a) and are vulnerable to adversarial manipulation (Tong et al., 2025). Self-assessment studies, such as Krumdick et al. (2025), demonstrate that while LLMs may show high overall agreement with human annotators, they often fail on subsets of tasks they cannot themselves solve. Similarly, Szymanski et al. (2025) found that LLM-based evaluation performs poorly in expert domains (e.g., dietetics, mental health), with substantially lower agreement than human judges. Beyond performance metrics, qualitative investigations raise broader concerns: Kapania et al. (2025) showed that when LLMs are used to simulate participants, their outputs often lack contextual depth, obscure marginalized perspectives, and amplify researcher positionality. There has been recent interest in using LLM-as-a-judge for evaluating education technologies or simulating student behavior (Dinucu-Jianu et al., 2025; Li et al., 2025b; Benedetto et al., 2024; Lu and Wang, 2024; Mannekote et al., 2025; Yuan et al., 2025; Zhang et al., 2024). However, educational evaluation is uniquely sensitive, as it involves assessing nuanced learner behaviors, metacognitive processes, and affective states—dimensions that the use of LLMs might oversimplify or misrepresent. Without stronger validation and safeguards, we advocate caution against relying on LLMs as judges to drive the design of tutoring systems, as premature adoption risks misaligning feedback with learners’ needs, and undermining trust in educational technologies.

### v2

@@ 5.3. Static vs Dynamic hints
We observed minimal differences in participant performance across the static and dynamic hint conditions, with average success of 7.61/10 and 7.88/10 questions respectively, compared to 6.36/10 questions in the no-hint control setting. Participants requested a total of 639 static hints, and 510 dynamic hints. Similar interaction patterns emerged across both hint conditions, illustrating substantial engagement with the system, without any design-imposed constraints (see Figure 7 top). Consequently, the adaptive mechanism in the dynamic condition was regularly engaged, enabling a valid comparison between static and adaptive hinting approaches.
(a) Distribution of prior actions succeeded by the “hint request” action.
(b) Average participant reported satisfaction ratings (1-5 Likert scale) grouped by the prior action.
(c) Percentage of hints yielding immediate success (y-axis) grouped by the prior action (x-axis).
Figure 7. Distribution of the prior actions succeeded by the “hint request” action (top), average reported satisfaction ratings (on a 1-5 Likert Scale) (middle), and the percentage of the hint request that led to immediate success after that hint grouped by prior actions (bottom). Orange (solid) bars denote the static hint generation statistics and blue (striped) bars denote the dynamic hint generation statistics; x-axis denotes the action taken prior to asking a thing, and y-axis denotes the statistic values. We observe that dynamic hints succeeding incorrect attempts achieve 31.32% immediate success, compared to 24.19% static hints, and obtain slightly higher satisfaction ratings ($\sim$0.13) illustrating that dynamic hints are able to adapt to learners’ responses to achieve higher answerability rate.
@@
The perceived difficulty of quiz was also rated similarly across all three sections, suggesting that neither hint strategy provided a decisive advantage at the level of overall outcomes. A Mann-Whitney U test (Mann and Whitney, 1947) revealed no statistically significant difference between the satisfaction ratings ($U=168351,p=0.365$) with negligible effect size (rank-biserial correlation = -0.033). Even after revealing the corresponding hint strategy to participants in the post-quiz survey, opinions remained divided. 20 out of 41 participants found static hints to be more helpful for answering the questions, while 23 out of 41 participants felt dynamic hints better improved their knowledge about the questions. Point-biserial correlation (Kornbrot, 2014) between post-section responses and post-quiz responses revealed that the reported quiz difficulty was not correlated with the post-disclosure preference ($corr=0.067,p=0.675$), indicating that the participants independently rated the quiz difficulty agnostic of hint helpfulness. The post-section hint helpfulness was moderately correlated with the post disclosure informativeness preference obtained from the posttest survey ($corr=0.471,p=0.002$), indicating that participants who found one hint strategy more helpful were likely to maintain their preference even after knowing the hint strategy. However, there is no correlation between the performance difference between static and dynamic, and the post-disclosure preference of answering assistance ($corr=0.063,p=0.694$), indicating that participants’ preferences were not influenced by their actual performance but rather by their subjective experience with the hints. Although neither strategy prevailed overall, participants described clear differences in how they experienced the two hinting approaches.
@@
Static hints–pre-generated in chains of four–were rigid in their progression and often provided diverse facts that enriched the context of the question. For example, ”Reflect on the types of landforms created by magma rising to the surface.”, or ”Research the type of radiation that includes visible light.” (see Figure 5). This pattern was reflected in reported informativeness scores: 39% of static hint chains were labeled informative, compared to only 35% of dynamic chains. However, Chi-squared test (Tallarida and Murray, 1987) revealed no statistically significant difference between the two strategies ($\chi^{2}=0.819,p=0.356$) with no association (Cramer’s V = 0.027). Providing a breadth of information allowed static hints to support sense-making and knowledge-building, but at times the progression felt misaligned with participants’ immediate needs.
@@
Dynamic hints, in contrast, were answer-fixated and more directly oriented toward guiding participants to the correct response. Because these hints were generated adaptively without awareness of the total number of hints in the chain, they lacked the deliberate pacing and planning of the static setting. However, this immediacy was effective when participants lacked sufficient prior knowledge to answer the question. For instance, hints like, ”Think about the rays named after a Greek letter”, or ”Consider a medication that starts with ”a” and is often used for headaches.” (see Figure 5). At the same time, the absence of chain-level planning made it harder to conceal answers: leakage was reported more often in dynamic hints (18%) than in static hints (15%). While Chi-squared test (Tallarida and Murray, 1987) revealed a statistically significance difference between the leakage responses ($\chi^{2}=5.35,p=0.021$), there was no association (Cramer’s V = 0.068). If the participant’s did not engage with the question, (i.e., not attempting any answers) the average satisfaction ratings for dynamic hints was worse than the static hints without engagement. While when they engaged with dynamic hints, their satisfaction ratings were comparable to static hints (see Figure 6 middle). While the participant reported satisfaction increased with increased engagement, the success rate of instances for dynamic instances were comparable to static hints (see Figure 6 middle). However, further investigation at individual hint level revealed that dynamic hints did adapt to incorrect responses. More specifically, the dynamic hints that succeeded an incorrect answer attempt were immediately followed by a successful answer for 31.32%, compared to 24.19% times for static hints (see Figure 7) bottom.
@@
Across participants, the overall sentiment remained consistent. Static hints were valued for their contextual richness, while dynamic hints were appreciated for their answer-directed efficiency. Minor exceptions arose due to randomized question distribution, with four participants noting one section to be easier than other, but the broader trend was stable. These findings suggest that static hints may be better suited for fostering understanding and providing transferable knowledge, whereas dynamic hints may be more effective in ensuring task success and engagement when learners are struggling.
@@ 5.4. Assessing effectiveness of automatic evaluation metrics
Automatic evaluation metrics are a key step toward scaling hint generation research: they allow us to compare systems without relying exclusively on human annotation. However, their usefulness depends on how well they capture the aspects of hint quality that learners themselves experience. To assess this alignment, we compared a set of automatic evaluation metrics against participant feedback collected in our human study. In total, 41 participants annotated 1,149 hints across 468 questions. Unlike our standardized automatic evaluation setup with a fixed chain size (see Appendix A), the hint chains here varied in length, providing a more realistic but also more heterogeneous basis for evaluation.
@@
We observed only weak correlations between information gain metrics and participants’ reported informativeness of hints for $InfoGain_{mean}$ ($corr_{pearson}=-0.121$) and $InfoGain_{comb}$ ($corr_{pearson}=-0.142$). These negative but shallow correlations highlight the limitations of relying on LLM-based measures of content quality, which often fail to capture the contextual insightfulness valued by participants. Although our study design did not directly test $Redundancy$ or $Consistency$, we also examined their correlations with reported satisfaction. $Redundancy$ showed a weak negative correlation ($corr_{pearson}=-0.245$), while $Consistency$ was essentially uncorrelated ($corr_{pearson}=-0.053$).
@@
For answer leakage, we were able to compare participant reports directly to our automatic metrics. The exact-match metric achieved high precision (0.834) but very low recall (0.075), indicating that it severely underestimated the degree of leakage present. Note that we implemented $leakage_{EM}$ with a simple string matching approach, that is unable to capture the nuances where hints reveal the answer in plural form (e.g., capillary vs capillaries), or change in part-of-speech (e.g., volcano vs volcanic). We also identified several indirect oversimplifications from hints that remain undetected by a simple syntactic overlap approach. For example, ”The bohr model works only for which atom?”, participants considered the hint ”It is the first element on the periodic table.” to have given away the answer, whereas it was undetected by the $leakage_{EM}$ metric. Conversely, the LLM-based leakage metric offered slightly higher recall (0.183) but suffered from low precision (0.266), indicating that while the learners didn’t consider certain hints to be oversimplifications, the evaluation LLM considered those hint chains to be so, overshooting the answer leakage predictions. For example, for the question ”What is the most abundant metal of the earth’s crust?”, and the provided hints ”Consider which metals are commonly found in rocks and soil.” and ”Think about the primary component of bauxite, a common ore.”, $leakage_{LLM}$ metric considers the hint chain to have leaked the answer (Aluminum), whereas the participant doesn’t indicate that in their responses. These results suggest that neither approach yet provides a reliable estimate of answer leakage from a learner’s perspective.
@@ 6. Discussion
Our findings highlight the complexity of automatic hint generation and the limitations of adopting off-the-shelf LLMs as tutoring systems. While participants consistently valued certain hint characteristics such as simplicity, novel knowledge perspectives, content diversity, and incremental scaffolding, their preferences diverged on various other aspects. For instance, some learners favored answer-focused hints that guided them toward the answer, while others found such feedback unhelpful in improving their learning. Beyond feedback on the hints generated by our system, participants also emphasized some unmet needs, with the hints unable to help bridge the linguistic barriers or connect well to their prior knowledge. In Section 6.1, we discuss the growing challenge of cognitive debt incurred from AI overreliance, and how automatic hint generation can serve as an avenue for addressing this problem. Drawing on our findings, we then discuss the design considerations for advancing hint generation beyond factual question answering in Section 6.2.
6.1. Hint Generation as a strategy against Cognitive Offloading
Recent years have seen rapid adoption of AI tools across everyday life including education, where students increasingly turn to AI assistance for help with assignments, writing, and problem-solving (Pitts et al., 2025a, b; Kanont et al., 2024; Lin, 2024). Prior work has documented this growing dependence, noting that learners often use these systems not only as aids but also as substitutes for their own cognitive effort (Adiyono et al., 2025). Recent studies further demonstrate the risks associated with excessive reliance on AI. For example, using electroencephalography signals, Kosmyna et al. (2025) found reduced neural coupling during AI-assisted essay writing. Similarly, utilizing a mixed-method approach over a large set of participants, Gerlich (2025) observed that heavier AI reliance correlates with weaker critical thinking scores among younger learners.
Our findings further contribute empirical evidence to this concern. Learners exhibited behaviors consistent with cognitive shortcuts, favoring hints that simplified answer retrieval rather than those that promoted reflection. This illustrates that over-reliance can arise even when systems are explicitly designed not to give full solutions, suggesting that interface-level decisions and scaffolding structure substantially influence learners’ help-seeking strategies.
While these risks are well-documented, recent systems-level approaches propose countermeasures such as systems that emphasize explainability (Hoq et al., 2025), promote metacognitive reflection (Kumar et al., 2024; Yuan and Hu, 2024), or deliberately withhold direct answers (Gupta et al., 2025). Our findings refine these design arguments; we show that simply withholding solutions is not sufficient as learners may still gravitate toward the most cognitively economical hint type. Instead, our data suggest that how hints are sequenced and when learners choose to request them are critical determinants of whether the system supports or undermines cognitive engagement.
When well-designed, hints can scaffold reasoning while preserving opportunities for productive struggle. Prior work shows that effective hints promote motivation and self-efficacy (Rogers et al., 2025; Mohammadi et al., 2025; VanLehn, 2011; Ma et al., 2014). Our findings reinforce this perspective. Participants generally rated hints as helpful for answering questions, with several explicitly noting their value in promoting deeper thinking. For instance, P-22 described hints as ”hints encouraged to make associations and engage in reflection from a scientific perspective,” P-37 noted they ”encouraged active thinking,” and P-40 highlighted they ”encouraged me to look for other ways to say things.” We found significant correlation between participants’ satisfaction with their information gain ($corr_{pearson}=0.4955$), indicating preference towards novel but relevant information conveyed through hints. At the same time, the contrast between static and adaptive hints in our study illustrates that subtle design choices can meaningfully shift the cognitive engagement patterns, and thereby the learning outcomes.
@@ 6.2. Design Implications Beyond Factual Hinting
The revised Bloom’s taxonomy of educational goals (Bloom et al., 1956; Anderson and Krathwohl, 2001) organizes cognitive processes from remember and understand through apply, analyze, evaluate, and create, while distinguishing knowledge goals between factual, conceptual, procedural, and metacognitive dimensions. Our study deliberately focused on factual knowledge at the lower levels of this hierarchy, where hints primarily supported remembering and understanding. This scope allowed us to systematically examine the role of hinting, while revealing several limitations in current technology that impede reliable implementation in authentic educational settings. For instance, challenges in robust answer assessment and misalignment between automatic evaluation metrics and learner perceptions. As one moves upward in Bloom’s hierarchy, such limitations become even more consequential: supporting conceptual reasoning, procedural mastery, or metacognitive reflection requires richer forms of adaptation, sensitivity to learner goals, and integration of affective and self-regulatory scaffolding. Here we outline design implications that extend beyond factual hinting, drawing on our findings to point toward future directions for building tutoring systems capable of supporting higher-order learning.
6.2.1. Building Trustworthy Foundations: Towards Reliable and Interpretable Tutoring Systems
Building AI tutoring systems that learners and educators can rely on requires more than producing fluent hints—it demands transparent, predictable, and pedagogically aligned foundations. Trustworthy systems make their reasoning traceable, provide consistent behavior across learners and contexts, and enable researchers to diagnose and mitigate errors. In the context of hint generation, this involves rigorously validating answer assessment for user interaction experience, building pedagogy-driven evaluation measures, and developing steerable and transparent hint generation systems that meet the learner’s goals. The following passages outline these challenges and potential pathways towards developing reliable and equitable AI tutors.
@@ 6.2.1. Building Trustworthy Foundations: Towards Reliable and Interpretable Tutoring Systems
Using LLMs for data synthesis and hint evaluation. LLMs are increasingly adopted in two major ways within educational technology research: generating synthetic data and serving as automated evaluators. For data synthesis, recent work shows that role-playing (Wang et al., 2025) and simulated tutoring dialogues can be used to fine-tune smaller models (Zhang et al., 2025) or optimize tutoring policies through LLM-based reward models (Dinucu-Jianu et al., 2025). These approaches promise scalability but also inherit the representational biases of the models generating the synthetic learners (Gautam et al., 2024; Ghosh et al., 2024), which may distort pedagogical signals, especially outside well-structured domains.
A parallel line of research uses LLMs as evaluators (“LLM-as-a-judge”) (Gu et al., 2024) to assess quality of educational content (Dinucu-Jianu et al., 2025; Li et al., 2025b; Benedetto et al., 2024; Lu and Wang, 2024; Mannekote et al., 2025; Qian et al., 2025). This method offers clear practical advantages over human studies with lower cost, fewer logistical barriers, and rapid iteration. However, educational evaluation is uniquely sensitive, as it involves assessing nuanced learner behaviors, metacognitive processes, and affective states—dimensions that the use of LLMs might oversimplify or misrepresent. Our own findings reflect this tension, the LLM-based metrics for informativeness and answer leakage showed weak alignment with learners’ judgments of hint quality, suggesting that LLM–based judges fail to capture the contextual insightfulness learners valued. Prior work similarly highlights that LLM evaluators overidentify student errors for math tutoring (Kakarla et al., 2024), and while fluent, often misaligned with instructor assessments for project report assessment (Dai et al., 2023).
@@
Investigations on LLM-based evaluators revealed that they are prone to systemic biases (Chen et al., 2024a; Ye et al., 2024; Shi et al., 2024; Wang et al., 2023b), and are vulnerable to adversarial manipulation (Tong et al., 2025). Self-assessment studies demonstrate that while LLMs may show high overall agreement with human annotators, they often fail on subsets of tasks they cannot themselves solve (Krumdick et al., 2025). Additionally, LLM-based evaluation performs poorly in expert domains (e.g., dietetics, mental health), with substantially lower agreement than human judges Krumdick et al. (2025). Beyond performance metrics, qualitative investigations raise broader concerns: Kapania et al. (2025) showed that when LLMs are used to simulate participants, their outputs often lack contextual depth, obscure marginalized perspectives, and amplify researcher positionality.
Together, these observations suggest that while LLM-driven synthesis and evaluation can help scale research, they are not yet dependable proxies for real learner experience. Progress will require metrics grounded in authentic interaction data and stronger validation protocols. Our annotated corpus of 1,100+ hints across 450+ questions provides one step toward more trustworthy evaluation frameworks, but future systems must adopt these techniques with caution to avoid reinforcing misaligned pedagogical signals or undermining trust in educational tools.
In-Context vs. Post-Training Approaches to Hint Generation. LLM-based hint generation generally follows two paths: (1) using in-context learning to produce feedback during a learner’s interaction, or (2) post-training models on task-specific corpora to develop specialized tutors. Because high-quality educational datasets are scarce, most prior work—including ours—follows the in-context approach, relying on off-the-shelf reasoning capabilities of LLMs to scaffold learners (Gabbay and Cohen, 2024; Dai et al., 2023). For instance, Pardos and Bhandari (2024) demonstrated that combining learner-authored prompts with chain-of-thought and self-consistency decoding (Wang et al., 2022) can improve algebra and statistics tutoring, though at higher computational cost. Similarly, Wan and Chen (2024) found that few-shot prompting enabled physics explanations that students preferred and domain experts judged as requiring only minimal edits.
Our findings align with these observations: in-context hints increased learner answerability by 13.8% relative to a no-hint control. However, our automatic evaluation over eighteen LLMs also shows substantial variance across model families and sizes–strong feedback is not a universal property of current LLMs (see Figure 15). This motivates attention to the second paradigm: targeted post-training. Prior work demonstrates its potential; for example, Scarlatos et al. (2025) trained a Llama3-8B (Grattafiori et al., 2024) model with direct preference optimization to outperform GPT models on mathematics tutoring benchmarks. Yet post-training is computationally demanding and hinges on carefully curated, pedagogically grounded datasets. As Dinucu-Jianu et al. (2025) showed, even within the same domain, different alignment strategies (e.g., supervised fine-tuning vs. RL-based optimization) can yield widely divergent results.
Taken together, these findings suggest that evaluating in-context learning is necessary but not sufficient for trustworthy educational systems. Reliable scaffolding will require progress on two fronts: building stronger benchmarks for model evaluation and developing principled post-training pipelines grounded in educational theory and interaction data.
Trade-offs between closed-source and open-source LLMs. A key distinction in our work is an explicit emphasis on open-source models, in contrast to much prior research that has relied on proprietary GPT systems (Jukiewicz, 2024). Closed-source models such as ChatGPT offer strong off-the-shelf performance, broad adoption among students, and rapid prototyping benefits for educational applications. However, their opacity limits opportunities for fine-grained analysis, reproducibility, and transparent evaluation of tutoring behavior Balloccu et al. (2024). Open-source alternatives, while often less capable without additional tuning, provide inspectability, controllability, and the ability to adapt models to curriculum-specific or domain-specific needs with modest compute budgets. Future work should consider the trade-offs between open-source and closed-source approaches: closed-source models remain valuable for advancing prototype performance, while open-source models are essential for reproducible research, equitable deployment, and sustained advances in educational AI.